An IV antibiotic has been ordered for a 5 year old admitted to the Peds unit...

  1. An IV antibiotic has been ordered for a 5 year old admitted to the Peds unit with cellulitis of the lower Right leg. Her weight is 42 pounds. The physician has ordered Clindamycin 40 mg/kg/day in 3 divided doses. You will be hanging the first dose. The child already has a primary IV of D5 ½ Normal Saline running at 60 ml/hour (maintenance rate).
  1. Calculate the dose in mgs expected for this child (show your calculation and round to the nearest hundreth):
Answer #1

Here baby's weight =42pounds

In kilograms =42/2.205

Total dose required =40mg/kg/day

ie, 40*19.47 =761.90mg/day

Child has to be given in 3 divided dose so, 761/3=253

The child required dose is 250md q8h